OpenStudy (anonymous):

Thank you, if you help. Solve the inequality. 2 + |t + 6| < 12 A. t > 4 or t < –11 B. t < 4 and t > 11 C. –16 < t < 4 D. –4 < t < 16

OpenStudy (anonymous):

\[\left| t+6 \right|+2<12\]\[\left| t+6 \right|<10\]when you want to solve these kinds of inequalities use this solution \[\left| x \right|<k \]\[-k <x <k\] so you can write -10<t+6<10 and you can easily solve it.